### v1

# BlockingPy: approximate nearest neighbours for blocking of records for entity resolution

Abstract
Entity resolution (probabilistic record linkage, deduplication) is a key step in scientific analysis and data science pipelines involving multiple data sources. The objective of entity resolution is to link records without identifiers that refer to the same entity (e.g., person, company). However, without identifiers, researchers need to specify which records to compare in order to calculate matching probability and reduce computational complexity. One solution is to deterministically block records based on some common variables, such as names, dates of birth or sex. However, this approach assumes that these variables are free of errors and completely observed, which is often not the case. To address this challenge, we have developed a Python package, BlockingPy, which utilises blocking via modern approximate nearest neighbour search and graph algorithms to significantly reduce the number of comparisons. In this paper, we present the design of the package, its functionalities and two case studies related to official statistics. We believe that the presented software will be useful for researchers (i.e., social scientists, economists or statisticians) interested in linking data from various sources.
keywords:
record linkage , deduplication , official statistics , economics
PACS:
89.20.Ff , 07.05.Kf
MSC:
68N01 , 68P10 , 68T10 , 68N01 , 62P25
journal: SoftwareX
Metadata
Nr. Code metadata description Metadata
C1 Current code version v0.1.14
C2 Permanent link to code/repository used for this code version https://github.com/ncn-foreigners/BlockingPy
C3 Permanent link to Reproducible Capsule https://blockingpy.readthedocs.io/en/latest/examples/record_linkage.html
C4 Legal Code License MIT License
C5 Code versioning system used Git
C6 Software code languages, tools, and services used Python
C7 Compilation requirements, operating environments & dependencies pandas, numpy, scipy, annoy, hnswlib, pynndescent, scikit-learn, networkx, nltk, voyager, faiss-cpu, mlpack
C8 If available Link to developer documentation/manual https://blockingpy.readthedocs.io/en/latest/index.html
C9 Support email for questions tymoteusz.strojny@gmail.com
Table 1: Code metadata (mandatory)
1 Motivation and significance
Entity resolution requires statistical and computational methodologies to accurately identify matching records across datasets without unique identifiers. This process underpins countless research endeavours across disciplines including epidemiology, economics, official statistics or historical research where data integration is fundamental to scientific discovery [1, 2]. Traditional blocking techniques, while computationally efficient, exhibit significant limitations when confronted with real-world data quality issues such as missing values, transliteration, typographical errors, and inconsistent formatting. These challenges, thoroughly documented by [3, 4, 5], have not been adequately addressed by existing software solutions, which often force researchers to choose between computational feasibility and linkage accuracy.
The scientific community has developed several tools to address record linkage, including R packages such as RecordLinkage [6], reclin2 [7], or fastLink [8], or Python implementations like splink [9], recordlinkage [10], Dedupe [11] or FEBRL [12]. However, these solutions typically rely on user-defined deterministic rules that may miss certain group of records or result in a large blocks of records when data irregularities are present (with the exception of the klsh, an R package that relies on the Locality Sensitive Hashing; [13]).
The BlockingPy package distinguishes itself by using state-of-the-art approximate nearest neighbour (ANN) search algorithms and graph-based indexing techniques that maintain robustness even with the imperfect data. This combination of modern algorithmic approaches represents a significant advancement over the fragmented workflow imposed by current tools, providing researchers with a seamless end-to-end pipeline that adaptively manages blocking strategies based on data characteristics. Users interact with BlockingPy through an user-friendly Python API that integrates with standard packages (e.g. pandas; cf. [14, 15]), allowing researchers to focus on analytical questions rather than technical implementation details of record linkage.
The structure of the paper is as follows. In Section 2 we briefly review the blocking procedures in existing Python packages. In Section 3 we provide details regarding the implementation ant the API. In Section 4 we provide two examples for probabilistic record linkage and deduplication. The paper finishes with brief review of the possible impact and conclusions.
2 Existing software covering blocking procedures
In this section, the focus is on Python, but readers are referred to the fastLink or reclin2 packages for more information on deterministic blocking procedures.
The splink package allows for deterministic blocking rules using function block_on() which specifies either the variables (e.g. block_on("city")) or more advanced combinations (e.g. block_on("substr(first_name, 1,1)", "surname")) which then are translated to SQL query for the DuckDB (via the BlockingRuleCreator class). More advanced users may be interested in creating their own rules which is possible through the CustomRule class. The splink allows users to specify the list of blocking rules to the Linker class. The splink allows users to inspect the blocking via the
splink.blocking_analysis method.
The recordlinkage package supports deterministic blocking via the recordlinkage.Index class and its methods. This package allows for deterministic blocking by BlockIndex() and SortedNeighbourhood() where specification of blocking variables should be provided by vector of column names and optionally window for the SortedNeighbourhood() which accounts for multiple mistakes.
The only software that provides similar functionalities is the blocklib package, which is part of the Anonlink project for privacy preserving record linkage (cf. [16, 17]). Unfortunately, at the time of writing, it has been two years since the last update, and this package no longer appears to be supported. We should also mention two recent packages: vicinity [18], which provides a similar API for the ANN algorithms described in Section 3, and semhash [19] that allows deduplication of records using the ANN algorithms via the vicinity. However, none of these packages provides an easy way to use this information for blocking of records for probabilistic record linkage and deduplication. Thus the BlockingPy provides a significant extensions to the existing Python packages.
3 Software description
3.1 Installation
BlockingPy is a Python package released under the MIT license. It can be installed from PyPI with popular package managers, e.g., pip or Poetry.
1pip install blockingpy
2poetry add blockingpy
3.2 Software functionalities
3.2.1 Blocking records
The core functionality of BlockingPy is to provide an efficient and scalable approach for blocking (also known as indexing) records in both record linkage and deduplication pipelines. To achieve this, users can input either previously computed Document-Term Matrices (DTMs) or raw text data, which the package transforms into DTMs by constructing character $n$-gram representations of the input.
Subsequently, similarity-based nearest neighbour search is being performed for each record in either the input dataset (deduplication) or query dataset (record linkage). This step is accomplished using one of the following Python implementations of ANN algorithms:
  • 1.
    Exact k-Nearest Neighbours (KNN) from mlpack and faiss packages.
  • 2.
    Locality Sensitive Hashing (LSH) from mlpack [20] and faiss packages.
  • 3.
    Hierarchical Navigable Small Worlds (HNSW) from faiss [21], hnswlib [22], and voyager [23] packages.
  • 4.
    Nearest Neighbour Descent (NND) from pynndescent [24] package.
  • 5.
    Random projections and NN trees from the annoy [25] package.
Users can fine-tune each algorithm according to their needs. Finally, BlockingPy groups records into blocks by identifying connected components in an undirected graph (using the igraph package; [26]), generated by previous neighbourhood search results. Figure 1 illustrates this workflow. By using the block() method with desired parameters on the Blocker instance, users can incorporate the above-mentioned process to obtain blocking results.
Figure 1: Blocking workflow in the BlockingPy package
3.2.2 Evaluation of blocking result
The blocking quality can be evaluated by providing ground truth blocks to either the block() or eval() methods. This allows users to assess the quality of the blocks computed by BlockingPy. The package implements batch processing for evaluation to allow processing of memory-intensive evaluations. The key evaluation information generated by the package is described in the sections below.
Reduction Ratio. Provides necessary details about the reduction in comparison pairs if the given blocks are applied to a further record linkage or deduplication procedure. For deduplication:
$\text{RR}_{\text{deduplication}}=1-\frac{\sum\limits_{i=1}^{k}\binom{|B_{i}|}{2}}{\binom{n}{2}},$
where $k$ is the total number of blocks, $n$ is the total number of records in the dataset, and $|B_{i}|$ is the number of records in the $i$-th block. $\sum\limits_{i=1}^{k}\binom{|B_{i}|}{2}$ is the number of comparisons after blocking, while $\binom{n}{2}$ is the total possible comparisons without blocking. For record linkage reduction rate is defined as follows
$\text{RR}_{\text{record\_linkage}}=1-\frac{\sum\limits_{i=1}^{k}|B_{i,x}|\cdot|B_{i,y}|}{(m\cdot n)},$
where $m$ and $n$ are the sizes of datasets $X$ and $Y$, and $k$ is the total number of blocks. The term $|B_{i,x}|$ is the number of unique records from dataset $X$ in the $i$-th block, while $|B_{i,y}|$ is the number of unique records from dataset $Y$ in the $i$-th block. The expression $\sum\limits_{i=1}^{k}|B_{i,x}|\cdot|B_{i,y}|$ is the number of comparisons after blocking.
Confusion Matrix. Presents results in comparison to ground-truth blocks in a pairwise manner (e.g., one true positive pair occurs when both records from the comparison pair belong to the same predicted block and to the same ground-truth block in the evaluation data frame.).
  • 1.
    True Positive (TP): Record pairs correctly matched in the same block.
  • 2.
    False Positive (FP): Records pairs identified as matches that are not true matches in the same block.
  • 3.
    True Negative (TN): Record pairs correctly identified as non-matches (different blocks)
  • 4.
    False Negative (FN): Records identified as non-matches that are true matches in the same block.
Evaluation metrics. Enables users to evaluate blocking across metrics as follows:
  • 1.
    Recall – $\frac{TP}{TP+FN}$.
  • 2.
    Precision – $\frac{TP}{TP+FP}$.
  • 3.
    Accuracy – $\frac{TP+TN}{TP+TN+FP+FN}$.
  • 4.
    Specificity – $\frac{TN}{TN+FP}$.
  • 5.
    F1 Score – $2\cdot\frac{\text{Precision}\times\text{Recall}}{\text{Precision}+\text{Recall}}$.
  • 6.
    False Positive Rate – $\frac{FP}{FP+TN}$.
  • 7.
    False Negative Rate – $\frac{FN}{FN+TP}$.
3.3 Software architecture
The architecture of BlockingPy provides a modular design, where each component is responsible for a specific aspect of the entire blocking workflow. This approach facilitates the integration of new ANN algorithms and modifications to existing functionality. The main components are illustrated in Figure 2 and their roles are described below.
Figure 2: The Architecture of the BlockingPy package
Blocker. The primary component of the BlockingPy architecture. It coordinates interaction between other components, managing the whole workflow of the package from handling data ingestion to generating results.
ANN Algorithms. Collection of ANN implementations with common interface for the straightforward addition of new algorithms. The parameters of each algorithm can be adjusted by providing the Controls dictionary to the control_ann parameter within the block method. An example of this is shown in Listing 1.
1blocker = Blocker()
2result = blocker.block(x=x, ann=’voyager’, control_ann=control_ann)
Listing 1: Setting the algorithm and fine-tuning with control_ann parameter
Input Processing. Manages the data pre-processing and transformation across multiple input formats. Converts dense and sparse matrices or raw text data to sparse data frames for efficient processing. Currently, we support transformation based on $n$-grams but in the future developments we plan to include support for vectors as in the semhash package.
Controls. Manages the configuration system for both text and ANN parameters. Includes customization for algorithm-specific options e.g., metric selection, search and query parameters, alongside tokenization rules and $n$-gram size. Listing 2 presents an example of Controls dictionary for the hnsw algorithm.
1control_ann = {
2 ’random_seed’: 2025
3 ’hnsw’: {
4 ’distance’: ’cosine’,
5 ’k_search’: 30,
6 ’n_threads’: 1,
7 ’path’: None,
8 ’M’: 25,
9 ’ef_c’: 200,
10 ’ef_s’: 200,
11 }
12}
Listing 2: Controls for hnsw algorithm
Input Validation. Ensures data and parameter correctness, while providing error messages for the user.
Blocking Result. Component responsible for management of the results and the corresponding information about the blocking process e.g., algorithm used, evaluation metrics, reduction ratio, or block distribution.
4 Illustrative examples
In the following section, we demonstrate BlockingPy package’s functionality on two examples of both record linkage and deduplication tasks. We aim to show users how user-friendly and effective the package is when blocking records is involved in entity resolution pipelines.
4.1 Record linkage
In this example, we use the Cis-Census datasets [27] to show the record linkage capabilities of BlockingPy. These datasets contain fictitious personal information, with the census dataset comprising 25,343 records and the cis (Customer Information System) dataset containing 24,613 records. Table 2 shows sample records from the census dataset after initial preparation conducted in Listing 3 for reference and the cis dataset is structured similarly.
NAME1 NAME2 SEX D M Y ENUMCAP ENUMPC
COUIE PRICE M 1 6 1960 1 WINDSOR ROAD DE03US
ABBIE PVICE F 9 11 1961 1 WINDSOR ROAD DE03US
LACEY PRICE F 7 2 1999 1 WINDSOR ROAD DE03US
SAMUEL PRICE M 13 4 1990 1 WINDSOR ROAD DE03US
JOSEPH PRICE M 20 4 1986 1 WINDSOR ROAD DE03US
Table 2: Example records from census dataset with variables used for txt column. Variables (changed here for better printing): NAME1 (first name), NAME2 (surname), SEX (gender), D, M, Y (birth date), ENUMCAP (address), ENUMPC (postal code).
1from blockingpy import Blocker
2import pandas as pd
3from blockingpy.datasets import load_census_cis_data
4
5census, cis = load_census_cis_data()
6
7census = census[["PERSON_ID","PERNAME1","PERNAME2","SEX","DOB_DAY","DOB_MON","DOB_YEAR","ENUMCAP","ENUMPC"]]
8cis = cis[["PERSON_ID","PERNAME1","PERNAME2","SEX", "DOB_DAY","DOB_MON","DOB_YEAR","ENUMCAP","ENUMPC"]]
Listing 3: Preprocessing the record linkage example data
Before blocking, we merge all fields of interest (without ID) into a single field (i.e., txt) that we will pass to the algorithm as shown in Listing 4. This step is performed for both datasets.
1for df in [census, cis]:
2 df[[’DOB_DAY’, ’DOB_MON’, ’DOB_YEAR’]] = (
3 df[[’DOB_DAY’, ’DOB_MON’, ’DOB_YEAR’]]
4 .astype("Int64")
5 .astype(str)
6 .replace(’<NA>’, ’’)
7 )
8 df.fillna(’’, inplace=True)
9
10 df[’txt’] = (
11 df[’PERNAME1’]
12 + df[’PERNAME2’]
13 + df[’SEX’]
14 + df[’DOB_DAY’]
15 + df[’DOB_MON’]
16 + df[’DOB_YEAR’]
17 + df[’ENUMCAP’]
18 + df[’ENUMPC’]
19 )
Listing 4: Creating the txt column
Subsequently, in Listing 5 we initialise Blocker and use the block() method and select the hnsw algorithm with default parameters to obtain the results. Tokenization and white-space removal are handled by the algorithm internally and can be modified via the control_txt parameter and the algorithm parameters can be fine-tuned via the control_ann argument.
1blocker = Blocker()
2rec_lin_result = blocker.block(
3 x=census[’txt’],
4 y=cis[’txt’],
5 ann=’hnsw’,
6 random_seed=42
7)
Listing 5: Record linkage blocking example
The output is a data frame containing rows with indices from both datasets (x, y) alongside the assigned block (an integer starting with 0) and distance computed with one of the available metrics. Listing 6 presents a snippet of the output in this example and Listing 7 shows the basic blocking information logged by the package where the first section informs about the algorithm used, the number of generated blocks, the number of features (created by shingling), the reduction ratio, and the second section displays the counts of blocks with each size, e.g., there were $591$ blocks created with each containing $3$ records.
1print(rec_lin_result.result.head())
2 x y block dist
30 17339 0 0 0.134151
41 9567 1 1 0.064307
52 10389 2 2 0.044183
63 24258 3 3 0.182125
74 3714 4 4 0.288487
Listing 6: Record linkage example of blocking result
1print(rec_lin_result)
2====================================================
3Blocking based on the hnsw method.
4Number of blocks: 23993
5Number of columns used for blocking: 1072
6Reduction ratio: 0.999961
7====================================================
8Distribution of the size of the blocks:
9Block Size | Number of Blocks
10 2 | 23388
11 3 | 591
12 4 | 13
13 5 | 1
Listing 7: Basic blocking information
We observe a reduction ratio of $0.9999$, which indicates that the possible number of candidate pairs is substantially reduced, thus lowering the computational complexity which is crucial for large-scale datasets.
With a unique personal identifier available in both datasets, we can evaluate the algorithm. For that, we create true_blocks data frame which represents indices from both datasets and their true block, as shown in Listing 8. In this case, every record in each block should point to the same unique entity. Furthermore, in Listing 9 we put the rec_lin_result object (from Listing 5) and the true_blocks in the eval method to obtain the evaluation results. In this example, we randomly sampled 1000 ground-truth pairs to showcase the evaluation functionality of the BlockingPy package.
1census[’x’] = range(len(census))
2cis[’y’] = range(len(cis))
3
4true_blocks = pd.merge(
5 left=census[[’PERSON_ID’, ’x’]],
6 right=cis[[’PERSON_ID’, ’y’]],
7 on=’PERSON_ID’
8)
9
10true_blocks[’block’] = range(len(true_blocks))
Listing 8: Preparation of true_blocks
1true_blocks = true_blocks.sample(1000, random_state=42)
2eval_result = blocker.eval(rec_lin_result, true_blocks[[’x’, ’y’, ’block’]])
Listing 9: Record Linkage evaluation code
Through this step, the user can access both the metrics and the confusion matrix. Evaluation is performed only for records that are available in true_blocks. We assume that we do not have knowledge about the records not included in the ground-truth data frame. Listing 10 and Listing 11 present the metrics and confusion matrix computed in this example.
1print(eval_result.metrics)
2recall 0.997000
3precision 1.000000
4fpr 0.000000
5fnr 0.003000
6accuracy 0.999997
7specificity 1.000000
8f1_score 0.998498
Listing 10: Evaluation Metrics
1print(eval_result.confusion)
2 Predicted Positive Predicted Negative
3Actual Positive 997 3
4Actual Negative 0 999000
Listing 11: Confusion Matrix
The results present BlockingPy’s ability to efficiently and accurately block records for record linkage tasks. The full version of this example is available on the documentation website.
4.2 Deduplication
In this example, our aim is to show how BlockingPy can be integrated into full entity resolution workflows. For that, we will present the deduplication of febrl1 dataset obtained from the recordlinkage package [10]. This dataset was generated using the FEBRL software [3] and contains 1000 records with fictitious personal information, of which 500 are original and 500 are duplicates with introduced errors. Firstly, we prepare the blocking key by imputing missing values with an empty string and concatenating fields with personal information into a single field, which we will use for blocking. After initialising Blocker and performing blocking on said data we are given the blocks by BlockingPy. Listing 12 presents this process, and the Listing 13 shows the outputted blocking information and the first five rows of the actual result.
1import recordlinkage
2from recordlinkage.datasets import load_febrl1
3from blockingpy import Blocker
4import pandas as pd
5import numpy as np
6
7df = load_febrl1()
8
9df = df.fillna(’’)
10df[’txt’] = df[’given_name’] + df[’surname’] + \
11 df[’street_number’] + df[’address_1’] + \
12 df[’address_2’] + df[’suburb’] + \
13 df[’postcode’] + df[’state’] + \
14 df[’date_of_birth’] + df[’soc_sec_id’]
15
16blocker = Blocker()
17blocking_result = blocker.block(
18 x=df[’txt’],
19 ann=’hnsw’,
20 random_seed=42
21)
Listing 12: Preprocessing and blocking
1print(blocking_result)
2========================================================
3Blocking based on the hnsw method.
4Number of blocks: 500
5Number of columns used for blocking: 1023
6Reduction ratio: 0.998999
7========================================================
8Distribution of the size of the blocks:
9Block Size | Number of Blocks
10 2 | 500
11
12print(blocking_result.result.head())
13 x y block dist
14 474 0 0 0.048375
15 330 1 1 0.038961
16 351 2 2 0.086690
17 290 3 3 0.024617
18 333 4 4 0.105662
Listing 13: Blocking Information
Both columns x and y refer to indices from the febrl1 alongside their designated block and distance measured between them with one of the available metrics (cosine in this example, the default for hnsw). Furthermore, the necessary step to make integration with recordlinkage possible is to add the block column to the original data frame, which we can do by transforming the blocking result as presented in Listing 14.
1result_df = blocking_result.result
2
3mapping_df = (
4 result_df
5 .melt(id_vars=[’block’], value_vars=[’x’, ’y’], value_name=’record_id’)
6 .drop_duplicates(subset=[’record_id’])
7)
8record_to_block = dict(zip(mapping_df[’record_id’], mapping_df[’block’]))
9new_data = df.copy()
10new_data[’block’] = [record_to_block.get(i) for i in range(len(df))]
Listing 14: Creating the block column
After calling the block method on the Index with the previously obtained block column in Listing 15 we can proceed with the usual recordlinkage workflow.
The full version of this example alongside the code can be found in the documentation.
1indexer = recordlinkage.Index()
2indexer.block(’block’)
3pairs = indexer.index(new_data)
Listing 15: Integration with recordlinkage
5 Impact
The proposed approach is used at Statistics Poland in the procedure for the linkage of records between administrative datasets as well as within one data set. For instance, we have used this approach to deduplicate the number of forced migrants crossing the Polish-Ukrainian border after the full-scale invasion of Ukraine by Russia on February 24, 2022. We also used this technique to determine the number of the residents of Ukraine under temporary protection in Poland as of March 31, 2023 [28]. The main motivation behind this was the problem of lack of identifiers as well as the transliteration from Ukrainian/Russian to Polish language (e.g. Олександр, should be Ołeksandr not Aleksandr).
In our opinion, this software opens up new research questions regarding optimal blocking strategies in administrative data integration, particularly in contexts where language differences and transliteration complexities exist. Researchers can now investigate how different blocking approaches affect linkage quality in multilingual settings, a question that was previously difficult to study systematically without appropriate tools. The software has significantly improved the pursuit of quality of official statistics by allowing for more accurate population estimates through better deduplication of records. Prior to this solution, Statistics Poland’s staff struggled with linking data without identifiers, applying user-defined rules which may have led to potential overestimation or underestimation of population flows.
Therefore, we believe that the proposed software will be of interest for researchers and practitioners who link data without identifiers or would like to apply large language models to detect duplicates but need to reduce costs by providing a small number of comparison buckets.
6 Conclusions
The current project is designed to be extensible so that it can accommodate new ANN algorithms and new input data formats. Any contributions are warmly welcome; see https://github.com/ncn-foreigners/BlockingPy/issues for a feature request and bug tracker. Future developments will focus on using semantic similarity as in the semhash package via custom encoders. Moreover, the framework can be extended to cover privacy preserving record linkage without directly sharing personally identifiable information.
Acknowledgements
This software is a part of Tymoteusz Strojny Bachelor’s thesis at Poznań University of Economics and Business, Poland. The BlockingPy is based on the blocking package in R (see https://ncn-foreigners.github.io/blocking/).
Work on this package is supported by the National Science Centre, OPUS 20 grant no. 2020/39/B/HS4/00941 (Towards census-like statistics for foreign-born populations – quality, data integration and estimation).
CRediT authorship contribution statement
Tymoteusz Strojny: Conceptualization, Software, Writing – Original Draft, Writing – Review & Editing; Maciej Beręsewicz: Conceptualization, Writing – Original Draft, Writing – Review & Editing, Supervision, Methodology, Project administration, Validation.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Declaration of generative AI and AI-assisted technologies in the writing process
During the preparation of this work the author(s) used DeepL and Claude.ai in order to proofread the paper. After using this the author(s) reviewed and edited the content as needed and take(s) full responsibility for the content of the published article.
Data availability
Data are publicly available.

### v2

@@
 Entity resolution (probabilistic record linkage, deduplication) is a key step in scientific analysis and data science pipelines involving multiple data sources.
-The objective of entity resolution is to link records without identifiers that refer to the same entity (e.g., person, company).
+The objective of entity resolution is to link records without common unique identifiers that refer to the same entity (e.g., person, company).
 However, without identifiers, researchers need to specify which records to compare in order to calculate matching probability and reduce computational complexity.
-One solution is to deterministically block records based on some common variables, such as names, dates of birth or sex.
+One solution is to deterministically block records based on some common variables, such as names, dates of birth or sex or use phonetic algorithms.
 However, this approach assumes that these variables are free of errors and completely observed, which is often not the case.
-To address this challenge, we have developed a Python package, BlockingPy, which utilises blocking via modern approximate nearest neighbour search and graph algorithms to significantly reduce the number of comparisons.
+To address this challenge, we have developed a Python package, BlockingPy, which uses blocking via modern approximate nearest neighbour search and graph algorithms to reduce the number of comparisons.
 In this paper, we present the design of the package, its functionalities and two case studies related to official statistics.
-We believe that the presented software will be useful for researchers (i.e., social scientists, economists or statisticians) interested in linking data from various sources.
+The presented software will be useful for researchers interested in linking data from various sources.
 keywords:
@@
 Nr. Code metadata description Metadata
-C1 Current code version v0.1.14
+C1 Current code version v0.1.15
 C2 Permanent link to code/repository used for this code version https://github.com/ncn-foreigners/BlockingPy
@@
 C6 Software code languages, tools, and services used Python
-C7 Compilation requirements, operating environments & dependencies pandas, numpy, scipy, annoy, hnswlib, pynndescent, scikit-learn, networkx, nltk, voyager, faiss-cpu, mlpack
+C7 Compilation requirements, operating environments & dependencies pandas, numpy, scipy, annoy, hnswlib, pynndescent, scikit-learn, networkx, nltk, voyager, faiss-cpu, mlpack, model2vec
 C8 If available Link to developer documentation/manual https://blockingpy.readthedocs.io/en/latest/index.html
@@
 Entity resolution requires statistical and computational methodologies to accurately identify matching records across datasets without unique identifiers.
-This process underpins countless research endeavours across disciplines including epidemiology, economics, official statistics or historical research where data integration is fundamental to scientific discovery [1, 2].
-Traditional blocking techniques, while computationally efficient, exhibit significant limitations when confronted with real-world data quality issues such as missing values, transliteration, typographical errors, and inconsistent formatting.
-These challenges, thoroughly documented by [3, 4, 5], have not been adequately addressed by existing software solutions, which often force researchers to choose between computational feasibility and linkage accuracy.
-The scientific community has developed several tools to address record linkage, including R packages such as RecordLinkage [6], reclin2 [7], or fastLink [8], or Python implementations like splink [9], recordlinkage [10], Dedupe [11] or FEBRL [12].
-However, these solutions typically rely on user-defined deterministic rules that may miss certain group of records or result in a large blocks of records when data irregularities are present (with the exception of the klsh, an R package that relies on the Locality Sensitive Hashing; [13]).
-The BlockingPy package distinguishes itself by using state-of-the-art approximate nearest neighbour (ANN) search algorithms and graph-based indexing techniques that maintain robustness even with the imperfect data.
-This combination of modern algorithmic approaches represents a significant advancement over the fragmented workflow imposed by current tools, providing researchers with a seamless end-to-end pipeline that adaptively manages blocking strategies based on data characteristics.
-Users interact with BlockingPy through an user-friendly Python API that integrates with standard packages (e.g.
-pandas; cf.
-[14, 15]), allowing researchers to focus on analytical questions rather than technical implementation details of record linkage.
+This process underpins countless research endeavours in disciplines such as epidemiology, economics, official statistics and historical research where data integration is fundamental to scientific discovery [1, 2].
+Traditional blocking techniques, while computationally efficient, exhibit limitations when confronted with various data quality issues such as missing values, transliteration, typographical errors, and inconsistent formatting.
+Existing deterministic blocking methods rely on user-defined exact-matching rules that assume key fields (e.g., name, date of birth) are error-free, potentially leading to missed links and false matches [3, 4, 5].
+Scientific community has developed several tools to address record linkage, including R packages such as RecordLinkage [6], reclin2 [7], fastLink [8], or Python implementations like splink [9], recordlinkage [10], Dedupe [11] or FEBRL [12].
+However, these solutions typically rely on user-defined deterministic rules that may miss certain groups of records or result in large blocks of records when data irregularities are present.
+Some solutions regarding these issues were proposed in the literature, for instance, the use of the SOUNDEX algorithm and other phonetic algorithms to block records on character strings that sound similar phonetically [13]; blocking using the Locality Sensitive Hashing (LSH; [14]), an approximate nearest neighbour (ANN) algorithm to search for similar records (see klsh package in R) or usage of Bloom filters, useful in particular for privacy-preserving record linkage (see, [15]; for implementation see blocklib in Python).
+The BlockingPy package distinguishes itself by using state-of-the-art ANN search algorithms and graph-based indexing techniques, which allow it to work even with imperfect data.
+BlockingPy implements an ANN-based blocking procedure ($n$-gram encoding or embeddings, neighbour search, graph components) behind a single pandas-style call, cutting custom rule-writing and setup time.
 The structure of the paper is as follows.
-In Section 2 we briefly review the blocking procedures in existing Python packages.
-In Section 3 we provide details regarding the implementation ant the API.
-In Section 4 we provide two examples for probabilistic record linkage and deduplication.
-The paper finishes with brief review of the possible impact and conclusions.
+In Section 2, we briefly review the blocking procedures in existing Python packages.
+In Section 3, we provide details regarding the implementation and the API.
+In Section 4, we provide two examples of probabilistic record linkage and deduplication.
+The paper finishes with a brief review of the possible impact and conclusions.
 2 Existing software covering blocking procedures
-In this section, the focus is on Python, but readers are referred to the fastLink or reclin2 packages for more information on deterministic blocking procedures.
-The splink package allows for deterministic blocking rules using function block_on() which specifies either the variables (e.g.
+In this section, the focus is on Python, but readers are referred to the fastLink and reclin2 packages for more information on deterministic blocking procedures.
+The splink package allows for deterministic blocking rules using the function block_on() which specifies either the variables (e.g.
 block_on("city")) or more advanced combinations (e.g.
-block_on("substr(first_name, 1,1)", "surname")) which then are translated to SQL query for the DuckDB (via the BlockingRuleCreator class).
-More advanced users may be interested in creating their own rules which is possible through the CustomRule class.
-The splink allows users to specify the list of blocking rules to the Linker class.
-The splink allows users to inspect the blocking via the
+block_on("substr(first_name, 1,1)", "surname")), which then are translated to a SQL query for the DuckDB (via the BlockingRuleCreator class).
+More advanced users may be interested in creating their own rules, which is possible through the CustomRule class.
+Splink allows users to specify the list of blocking rules for the Linker class.
+Splink allows users to inspect the blocking via the
 splink.blocking_analysis method.
-The recordlinkage package supports deterministic blocking via the recordlinkage.Index class and its methods.
-This package allows for deterministic blocking by BlockIndex() and SortedNeighbourhood() where specification of blocking variables should be provided by vector of column names and optionally window for the SortedNeighbourhood() which accounts for multiple mistakes.
-The only software that provides similar functionalities is the blocklib package, which is part of the Anonlink project for privacy preserving record linkage (cf.
+Additionally, splink supports blocking using phonetic algorithms such as Soundex, Metaphone, or Double Metaphone either by applying them directly within the blocking rules (e.g., block_on("soundex(surname)")), or by preprocessing columns to store phonetic encodings separately.
+Recordlinkage package supports deterministic blocking via the recordlinkage.Index class and its methods.
+This package allows for deterministic blocking using BlockIndex() and SortedNeighbourhood() where specification of blocking variables should be provided by a vector of column names and optionally window for the SortedNeighbourhood(), which accounts for multiple mistakes.
+The only software that provides similar functionality is the blocklib package, which is part of the Anonlink project for privacy-preserving record linkage (cf.
 [16, 17]).
-Unfortunately, at the time of writing, it has been two years since the last update, and this package no longer appears to be supported.
-We should also mention two recent packages: vicinity [18], which provides a similar API for the ANN algorithms described in Section 3, and semhash [19] that allows deduplication of records using the ANN algorithms via the vicinity.
-However, none of these packages provides an easy way to use this information for blocking of records for probabilistic record linkage and deduplication.
-Thus the BlockingPy provides a significant extensions to the existing Python packages.
+We should also mention three recent packages: vicinity [18], which provides a similar API for the ANN algorithms described in Section 3, semhash [19], which allows deduplication of records using the ANN algorithms via the vicinity and LinkTransformer [20], which uses kNN-based embedding comparisons and offers exact-match blocking.
+However, none of these packages provide a way to use this information for blocking of records for probabilistic record linkage and deduplication.
+Thus, BlockingPy provides an extension to the existing Python package ecosystem.
 3 Software description
 3.1 Installation
-BlockingPy is a Python package released under the MIT license. It can be installed from PyPI with popular package managers, e.g., pip or Poetry.
+BlockingPy is a Python package released under the MIT license. It can be installed from PyPI using package managers such as pip or Poetry.
 1pip install blockingpy
@@
 3.2 Software functionalities
+This subsection describes the main functionalities provided by the BlockingPy package.
 3.2.1 Blocking records
@@
 To achieve this, users can input either previously computed Document-Term Matrices (DTMs) or raw text data, which the package transforms into DTMs by constructing character $n$-gram representations of the input.
-Subsequently, similarity-based nearest neighbour search is being performed for each record in either the input dataset (deduplication) or query dataset (record linkage).
+Alternatively, users can choose to represent the input data as vector embeddings.
+Subsequently, similarity-based nearest neighbour search is performed for each record in either the input dataset (deduplication) or query dataset (record linkage).
 This step is accomplished using one of the following Python implementations of ANN algorithms:
@@
 Users can fine-tune each algorithm according to their needs.
-Finally, BlockingPy groups records into blocks by identifying connected components in an undirected graph (using the igraph package; [26]), generated by previous neighbourhood search results.
-Figure 1 illustrates this workflow.
-By using the block() method with desired parameters on the Blocker instance, users can incorporate the above-mentioned process to obtain blocking results.
-Figure 1: Blocking workflow in the BlockingPy package
+Finally, BlockingPy groups records into blocks by identifying connected components in an undirected graph (using the igraph package [27]), generated by previous neighbourhood search results.
+By using the block() method with the desired parameters on the Blocker instance, users can run the above-mentioned process to obtain blocking results.
 3.2.2 Evaluation of blocking result
@@
 This allows users to assess the quality of the blocks computed by BlockingPy.
-The package implements batch processing for evaluation to allow processing of memory-intensive evaluations.
 The key evaluation information generated by the package is described in the sections below.
@@
 where $k$ is the total number of blocks, $n$ is the total number of records in the dataset, and $|B_{i}|$ is the number of records in the $i$-th block.
-$\sum\limits_{i=1}^{k}\binom{|B_{i}|}{2}$ is the number of comparisons after blocking, while $\binom{n}{2}$ is the total possible comparisons without blocking.
-For record linkage reduction rate is defined as follows
+$\sum\limits_{i=1}^{k}\binom{|B_{i}|}{2}$ is the number of comparisons after blocking, while $\binom{n}{2}$ is the total number of possible comparisons without blocking.
+For record linkage the reduction ratio is defined as follows
 $\text{RR}_{\text{record\_linkage}}=1-\frac{\sum\limits_{i=1}^{k}|B_{i,x}|\cdot|B_{i,y}|}{(m\cdot n)},$
@@
 The expression $\sum\limits_{i=1}^{k}|B_{i,x}|\cdot|B_{i,y}|$ is the number of comparisons after blocking.
-Confusion Matrix.
-Presents results in comparison to ground-truth blocks in a pairwise manner (e.g., one true positive pair occurs when both records from the comparison pair belong to the same predicted block and to the same ground-truth block in the evaluation data frame.).
+Confusion matrix presents results in comparison to ground-truth blocks in a pairwise manner (e.g., one true positive pair occurs when both records from the comparison pair belong to the same predicted block and to the same ground-truth block in the evaluation data frame).
   • 1.
@@
     False Negative (FN): Records identified as non-matches that are true matches in the same block.
-Evaluation metrics. Enables users to evaluate blocking across metrics as follows:
-  • 1.
-    Recall – $\frac{TP}{TP+FN}$.
-  • 2.
-    Precision – $\frac{TP}{TP+FP}$.
-  • 3.
-    Accuracy – $\frac{TP+TN}{TP+TN+FP+FN}$.
-  • 4.
-    Specificity – $\frac{TN}{TN+FP}$.
-  • 5.
-    F1 Score – $2\cdot\frac{\text{Precision}\times\text{Recall}}{\text{Precision}+\text{Recall}}$.
-  • 6.
-    False Positive Rate – $\frac{FP}{FP+TN}$.
-  • 7.
-    False Negative Rate – $\frac{FN}{FN+TP}$.
+Finally, table 2 presents metrics that enables users to evaluate blocking procedures.
+Metric Formula Metric Formula
+Recall $\frac{TP}{TP+FN}$ Accuracy $\frac{TP+TN}{TP+TN+FP+FN}$
+Precision $\frac{TP}{TP+FP}$ Specificity $\frac{TN}{TN+FP}$
+F1 Score $2\cdot\frac{\text{Precision}\times\text{Recall}}{\text{Precision}+\text{Recall}}$ False Positive Rate $\frac{FP}{FP+TN}$
+False Negative Rate $\frac{FN}{FN+TP}$
+Table 2: Evaluation Metrics
 3.3 Software architecture
@@
 This approach facilitates the integration of new ANN algorithms and modifications to existing functionality.
-The main components are illustrated in Figure 2 and their roles are described below.
+The main components are illustrated in Figure 1, and their roles are described below.
 Figure 1: The Architecture of the BlockingPy package
@@
 The primary component of the BlockingPy architecture.
-It coordinates interaction between other components, managing the whole workflow of the package from handling data ingestion to generating results.
+It coordinates interactions between other components, managing the whole workflow of the package from handling data ingestion to generating results.
 ANN Algorithms.
-Collection of ANN implementations with common interface for the straightforward addition of new algorithms.
+Collection of ANN implementations with a common interface for the straightforward addition of new algorithms.
 The parameters of each algorithm can be adjusted by providing the Controls dictionary to the control_ann parameter within the block method.
@@
 Manages the data pre-processing and transformation across multiple input formats.
-Converts dense and sparse matrices or raw text data to sparse data frames for efficient processing.
-Currently, we support transformation based on $n$-grams but in the future developments we plan to include support for vectors as in the semhash package.
+Converts dense matrices, sparse matrices, and raw text data to sparse data frames for efficient processing.
+Currently, we support transformation based on $n$-grams and vector embeddings.
 Controls.
 Manages the configuration system for both text and ANN parameters.
-Includes customization for algorithm-specific options e.g., metric selection, search and query parameters, alongside tokenization rules and $n$-gram size.
-Listing 2 presents an example of Controls dictionary for the hnsw algorithm.
+Includes customization for algorithm-specific options e.g., metric selection, search and query parameters, alongside tokenization rules and $n$-gram size, or alternatively the choice of an embedding model.
+Listing 2 presents an example of a Controls dictionary for the hnsw algorithm.
 1control_ann = {
@@
 Listing 2: Controls for hnsw algorithm
-Input Validation. Ensures data and parameter correctness, while providing error messages for the user.
+Input Validation. Ensures validity of the data and parameters while providing error messages for the user.
 Blocking Result.
-Component responsible for management of the results and the corresponding information about the blocking process e.g., algorithm used, evaluation metrics, reduction ratio, or block distribution.
+The component responsible for managing results and the corresponding information about the blocking process e.g., algorithm used, evaluation metrics, reduction ratio, and block distribution.
 4 Illustrative examples
-In the following section, we demonstrate BlockingPy package’s functionality on two examples of both record linkage and deduplication tasks.
-We aim to show users how user-friendly and effective the package is when blocking records is involved in entity resolution pipelines.
+In the following section, we demonstrate the BlockingPy package’s functionality using two examples: one for record linkage and one for deduplication.
+We aim to show the ease of use and effectiveness of the package in blocking records for entity resolution.
 4.1 Record linkage
-In this example, we use the Cis-Census datasets [27] to show the record linkage capabilities of BlockingPy.
+In this example, we use the Cis-Census datasets (for more information see [28]) to show the record linkage capabilities of BlockingPy.
 These datasets contain fictitious personal information, with the census dataset comprising 25,343 records and the cis (Customer Information System) dataset containing 24,613 records.
-Table 2 shows sample records from the census dataset after initial preparation conducted in Listing 3 for reference and the cis dataset is structured similarly.
+Table 3 shows sample records from the census dataset after the initial preparation described in Listing 3.
+The cis dataset is structured similarly.
 NAME1 NAME2 SEX D M Y ENUMCAP ENUMPC
@@
 Listing 4: Creating the txt column
-Subsequently, in Listing 5 we initialise Blocker and use the block() method and select the hnsw algorithm with default parameters to obtain the results.
-Tokenization and white-space removal are handled by the algorithm internally and can be modified via the control_txt parameter and the algorithm parameters can be fine-tuned via the control_ann argument.
+Subsequently, in Listing 5 we initialise Blocker, invoke the block() method, and select the hnsw algorithm with default parameters to obtain the results.
+The tokenization and white-space removal are handled by the algorithm internally and can be modified via the control_txt parameter.
+Additional algorithm-specific parameters can be fine-tuned via the control_ann argument.
 1blocker = Blocker()
@@
 The output is a data frame containing rows with indices from both datasets (x, y) alongside the assigned block (an integer starting with 0) and distance computed with one of the available metrics.
-Listing 6 presents a snippet of the output in this example and Listing 7 shows the basic blocking information logged by the package where the first section informs about the algorithm used, the number of generated blocks, the number of features (created by shingling), the reduction ratio, and the second section displays the counts of blocks with each size, e.g., there were $591$ blocks created with each containing $3$ records.
+Listing 6 presents a snippet of the output in this example.
+Listing 7 shows the basic blocking information logged by the package.
+The first section provides details about the algorithm used, the number of generated blocks, the number of columns used for blocking (created by shingling, a preprocessing step that converts strings into overlapping letter groups used as input to the ANN algorithm), and the reduction ratio.
+The second section displays the distribution of block sizes; for example, there were $591$ blocks created, each containing $3$ records.
 1print(rec_lin_result.result.head())
@@
 4Number of blocks: 23993
-5Number of columns used for blocking: 1072
+5Number of columns created for blocking: 1072
 6Reduction ratio: 0.999961
@@
 Listing 7: Basic blocking information
-We observe a reduction ratio of $0.9999$, which indicates that the possible number of candidate pairs is substantially reduced, thus lowering the computational complexity which is crucial for large-scale datasets.
+We observe a reduction ratio of $0.9999$, which indicates that the possible number of candidate pairs is substantially reduced, thus lowering the computational complexity, which is crucial for large-scale datasets.
 With a unique personal identifier available in both datasets, we can evaluate the algorithm.
-For that, we create true_blocks data frame which represents indices from both datasets and their true block, as shown in Listing 8.
+For that, we create a true_blocks data frame which represents indices from both datasets and their true block, as shown in Listing 8.
 In this case, every record in each block should point to the same unique entity.
-Furthermore, in Listing 9 we put the rec_lin_result object (from Listing 5) and the true_blocks in the eval method to obtain the evaluation results.
+Furthermore, in Listing 9 we pass the rec_lin_result object (from Listing 5) and the true_blocks to the eval method to obtain the evaluation results.
 In this example, we randomly sampled 1000 ground-truth pairs to showcase the evaluation functionality of the BlockingPy package.
@@
 Listing 9: Record Linkage evaluation code
-Through this step, the user can access both the metrics and the confusion matrix.
+Through this step, the user can access both the evaluation metrics and the confusion matrix.
 Evaluation is performed only for records that are available in true_blocks.
-We assume that we do not have knowledge about the records not included in the ground-truth data frame.
+We assume that we have no knowledge about the records not included in the ground-truth data frame.
 Listing 10 and Listing 11 present the metrics and confusion matrix computed in this example.
@@
 In this example, our aim is to show how BlockingPy can be integrated into full entity resolution workflows.
-For that, we will present the deduplication of febrl1 dataset obtained from the recordlinkage package [10].
-This dataset was generated using the FEBRL software [3] and contains 1000 records with fictitious personal information, of which 500 are original and 500 are duplicates with introduced errors.
+For that, we will present the deduplication of the febrl1 dataset obtained from the recordlinkage package [10].
+This dataset was generated using the FEBRL software and contains 1000 records with fictitious personal information, of which 500 are original and 500 are duplicates with introduced errors.
 Firstly, we prepare the blocking key by imputing missing values with an empty string and concatenating fields with personal information into a single field, which we will use for blocking.
-After initialising Blocker and performing blocking on said data we are given the blocks by BlockingPy.
-Listing 12 presents this process, and the Listing 13 shows the outputted blocking information and the first five rows of the actual result.
+After initialising Blocker and performing blocking on said data, we are given the blocks by BlockingPy.
+Listing 12 presents this process, and the Listing 13 shows the blocking information output and the first five rows of the actual result.
 1import recordlinkage
@@
 Listing 13: Blocking Information
-Both columns x and y refer to indices from the febrl1 alongside their designated block and distance measured between them with one of the available metrics (cosine in this example, the default for hnsw).
-Furthermore, the necessary step to make integration with recordlinkage possible is to add the block column to the original data frame, which we can do by transforming the blocking result as presented in Listing 14.
-1result_df = blocking_result.result
-2
-3mapping_df = (
-4 result_df
-5 .melt(id_vars=[’block’], value_vars=[’x’, ’y’], value_name=’record_id’)
-6 .drop_duplicates(subset=[’record_id’])
-7)
-8record_to_block = dict(zip(mapping_df[’record_id’], mapping_df[’block’]))
-9new_data = df.copy()
-10new_data[’block’] = [record_to_block.get(i) for i in range(len(df))]
+Both columns x and y refer to indices from the febrl1 dataset alongside their designated block and distance measured between them using one of the available metrics (cosine in this example, the default for hnsw).
+Furthermore, the necessary step to make integration with recordlinkage possible is to add the block column to the original data frame, which we can do with the add_block_column method as presented in Listing 14.
+1df[’id’] = range(len(df))
+2df_final = blocking_result.add_block_column(df, id_col_left=’id’)
 Listing 14: Creating the block column
 After calling the block method on the Index with the previously obtained block column in Listing 15 we can proceed with the usual recordlinkage workflow.
-The full version of this example alongside the code can be found in the documentation.
+The full version of this example and many others, alongside the code, can be found in the documentation.
 1indexer = recordlinkage.Index()
 2indexer.block(’block’)
-3pairs = indexer.index(new_data)
+3pairs = indexer.index(df_final)
 Listing 15: Integration with recordlinkage
@@
 For instance, we have used this approach to deduplicate the number of forced migrants crossing the Polish-Ukrainian border after the full-scale invasion of Ukraine by Russia on February 24, 2022.
-We also used this technique to determine the number of the residents of Ukraine under temporary protection in Poland as of March 31, 2023 [28].
+We also used this technique to determine the number of residents of Ukraine under temporary protection in Poland as of March 31, 2023 [29].
 The main motivation behind this was the problem of lack of identifiers as well as the transliteration from Ukrainian/Russian to Polish language (e.g.
 Олександр, should be Ołeksandr not Aleksandr).
-In our opinion, this software opens up new research questions regarding optimal blocking strategies in administrative data integration, particularly in contexts where language differences and transliteration complexities exist.
+This software opens up new research questions regarding optimal blocking strategies in administrative data integration, particularly in contexts where language differences and transliteration complexities exist.
 Researchers can now investigate how different blocking approaches affect linkage quality in multilingual settings, a question that was previously difficult to study systematically without appropriate tools.
-The software has significantly improved the pursuit of quality of official statistics by allowing for more accurate population estimates through better deduplication of records.
-Prior to this solution, Statistics Poland’s staff struggled with linking data without identifiers, applying user-defined rules which may have led to potential overestimation or underestimation of population flows.
-Therefore, we believe that the proposed software will be of interest for researchers and practitioners who link data without identifiers or would like to apply large language models to detect duplicates but need to reduce costs by providing a small number of comparison buckets.
+The software has meaningfully improved the pursuit of quality of official statistics by allowing for more accurate population estimates through better deduplication of records.
+Before implementing this solution, Statistics Poland relied on user-defined rules for linking data without identifiers, which often resulted in imprecise population flow estimates due to missed matches or false positives.
+We believe that the proposed software will be of interest to researchers and practitioners who link data without identifiers or would like to apply large language models to detect duplicates but need to reduce costs by providing a small number of comparison buckets.
 6 Conclusions
@@
 Any contributions are warmly welcome; see https://github.com/ncn-foreigners/BlockingPy/issues for a feature request and bug tracker.
-Future developments will focus on using semantic similarity as in the semhash package via custom encoders.
-Moreover, the framework can be extended to cover privacy preserving record linkage without directly sharing personally identifiable information.
+Future updates will focus on adding functionality for privacy-preserving record linkage without directly sharing personally identifiable information.
 Acknowledgements
-This software is a part of Tymoteusz Strojny Bachelor’s thesis at Poznań University of Economics and Business, Poland.
-The BlockingPy is based on the blocking package in R (see https://ncn-foreigners.github.io/blocking/).
+We thank the reviewers for their comments on this manuscript.
+This software is a part of Tymoteusz Strojny’s Bachelor’s thesis at Poznań University of Economics and Business, Poland.
+BlockingPy is based on the blocking package in R (see https://ncn-foreigners.github.io/blocking/).
 Work on this package is supported by the National Science Centre, OPUS 20 grant no.
@@
 Declaration of generative AI and AI-assisted technologies in the writing process
-During the preparation of this work the author(s) used DeepL and Claude.ai in order to proofread the paper.
+During the preparation of this work the author(s) used DeepL and Claude.ai to proofread the paper.
 After using this the author(s) reviewed and edited the content as needed and take(s) full responsibility for the content of the published article.
 Data availability
-Data are publicly available.
+The datasets used in Section 4 are included with the BlockingPy package and are also publicly available from their original sources.

### v3

@@
 However, this approach assumes that these variables are free of errors and completely observed, which is often not the case.
-To address this challenge, we have developed a Python package, BlockingPy, which uses blocking via modern approximate nearest neighbour search and graph algorithms to reduce the number of comparisons.
+To address this challenge, we have developed a Python package, BlockingPy, which uses blocking using modern approximate nearest neighbour search and graph algorithms to reduce the number of comparisons.
+The package supports both CPU and GPU execution.
 In this paper, we present the design of the package, its functionalities and two case studies related to official statistics.
@@
 Nr. Code metadata description Metadata
-C1 Current code version v0.1.15
+C1 Current code version v0.2.8
 C2 Permanent link to code/repository used for this code version https://github.com/ncn-foreigners/BlockingPy
@@
 C6 Software code languages, tools, and services used Python
-C7 Compilation requirements, operating environments & dependencies pandas, numpy, scipy, annoy, hnswlib, pynndescent, scikit-learn, networkx, nltk, voyager, faiss-cpu, mlpack, model2vec
+C7 Compilation requirements, operating environments & dependencies pandas, numpy, scipy, annoy, hnswlib, pynndescent, scikit-learn, igraph, nltk, voyager, faiss-cpu, faiss-gpu-cuvs, mlpack, model2vec
 C8 If available Link to developer documentation/manual https://blockingpy.readthedocs.io/en/latest/index.html
@@
 Some solutions regarding these issues were proposed in the literature, for instance, the use of the SOUNDEX algorithm and other phonetic algorithms to block records on character strings that sound similar phonetically [13]; blocking using the Locality Sensitive Hashing (LSH; [14]), an approximate nearest neighbour (ANN) algorithm to search for similar records (see klsh package in R) or usage of Bloom filters, useful in particular for privacy-preserving record linkage (see, [15]; for implementation see blocklib in Python).
-The BlockingPy package distinguishes itself by using state-of-the-art ANN search algorithms and graph-based indexing techniques, which allow it to work even with imperfect data.
+The BlockingPy package uses state-of-the-art ANN search algorithms and graph-based indexing techniques, which allow it to work even with imperfect data.
 BlockingPy implements an ANN-based blocking procedure ($n$-gram encoding or embeddings, neighbour search, graph components) behind a single pandas-style call, cutting custom rule-writing and setup time.
@@
 Additionally, splink supports blocking using phonetic algorithms such as Soundex, Metaphone, or Double Metaphone either by applying them directly within the blocking rules (e.g., block_on("soundex(surname)")), or by preprocessing columns to store phonetic encodings separately.
+In online documentation we show that using splink with BlockingPy together results in significantly improvement in the performance metrics by capturing comparison pairs that would otherwise be missed.
 Recordlinkage package supports deterministic blocking via the recordlinkage.Index class and its methods.
 This package allows for deterministic blocking using BlockIndex() and SortedNeighbourhood() where specification of blocking variables should be provided by a vector of column names and optionally window for the SortedNeighbourhood(), which accounts for multiple mistakes.
-The only software that provides similar functionality is the blocklib package, which is part of the Anonlink project for privacy-preserving record linkage (cf.
-[16, 17]).
-We should also mention three recent packages: vicinity [18], which provides a similar API for the ANN algorithms described in Section 3, semhash [19], which allows deduplication of records using the ANN algorithms via the vicinity and LinkTransformer [20], which uses kNN-based embedding comparisons and offers exact-match blocking.
-However, none of these packages provide a way to use this information for blocking of records for probabilistic record linkage and deduplication.
-Thus, BlockingPy provides an extension to the existing Python package ecosystem.
+The only software that provides similar functionality is the blocklib package (cf.
+[16, 17]) and vicinity [18].
+The first uses probability signature and LSH-based Lambda Fold Redundant as described in [16].
+The second uses semhash [19], which allows deduplication of records using kNN-based embedding comparisons and offers exact-match blocking.
+In contrast, BlockingPy allows various inputs as well as ANN algorithms and graphs, focusing on providing accurate blocking for probabilistic record linkage and deduplication.
+In Appendix Appendix A.
+Numerical comparison with blocklib, we compare BlockingPy with blocklib in terms of computational efficiency, accuracy, and number of pairs from the blocking procedures.
+While blocklib is faster, it results in a significantly higher number of missed pairs and several orders of magnitude more pairs.
+We also provide comparison with the GPU version of the package.
 3 Software description
@@
 2poetry add blockingpy
+GPU option. We also provide blockingpy-gpu with CUDA-enabled GPU execution:
+1# 1) Create env
+2mamba create -n blockingpy-gpu python=3.10 -y
+3conda activate blockingpy-gpu
+4
+5# 2) Install FAISS GPU (tested cuVS nightly)
+6mamba install -c pytorch/label/nightly \
+7 faiss-gpu-cuvs=1.11.0=py3.10_ha3bacd1_55_cuda12.4.0_nightly -y
+8
+9# 3) Install BlockingPy (GPU build)
+10pip install blockingpy-gpu
+Notes: Import remains import blockingpy.
 3.2 Software functionalities
@@
 The core functionality of BlockingPy is to provide an efficient and scalable approach for blocking (also known as indexing) records in both record linkage and deduplication pipelines.
-To achieve this, users can input either previously computed Document-Term Matrices (DTMs) or raw text data, which the package transforms into DTMs by constructing character $n$-gram representations of the input.
+Scalability is achieved mainly due to the ANN libraries that we use in our implementations.
+Users can input either previously computed Document-Term Matrices (DTMs) or raw text data, which the package transforms into DTMs by constructing character $n$-gram representations of the input.
 Alternatively, users can choose to represent the input data as vector embeddings.
@@
     Random projections and NN trees from the annoy [25] package.
+  • 6.
+    Inverted File Flat (IVF), IVF with Product Quantization (IVFPQ) and CAGRA [26] from faiss (GPU) package.
 Users can fine-tune each algorithm according to their needs.
@@
 ANN Algorithms.
-Collection of ANN implementations with a common interface for the straightforward addition of new algorithms.
+This is a collection of ANN implementations with a common interface for the straightforward addition of new algorithms.
 The parameters of each algorithm can be adjusted by providing the Controls dictionary to the control_ann parameter within the block method.
 An example of this is shown in Listing 1.
+All ANN backends share one interface across CPU and GPU.
 1blocker = Blocker()
@@
 Manages the data pre-processing and transformation across multiple input formats.
-Converts dense matrices, sparse matrices, and raw text data to sparse data frames for efficient processing.
+Converts dense matrices, sparse matrices, and raw text data to Numpy arrays or Scipy CSR matrices via the DataHandler class for efficient processing.
 Currently, we support transformation based on $n$-grams and vector embeddings.
@@
 The proposed approach is used at Statistics Poland in the procedure for the linkage of records between administrative datasets as well as within one dataset.
-For instance, we have used this approach to deduplicate the number of forced migrants crossing the Polish-Ukrainian border after the full-scale invasion of Ukraine by Russia on February 24, 2022.
+For instance, we have used this approach to deduplicate records of forced migrants crossing the Polish-Ukrainian border after the full-scale invasion of Ukraine by Russia on February 24, 2022.
 We also used this technique to determine the number of residents of Ukraine under temporary protection in Poland as of March 31, 2023 [29].
@@
 Олександр, should be Ołeksandr not Aleksandr).
-This software opens up new research questions regarding optimal blocking strategies in administrative data integration, particularly in contexts where language differences and transliteration complexities exist.
-Researchers can now investigate how different blocking approaches affect linkage quality in multilingual settings, a question that was previously difficult to study systematically without appropriate tools.
-The software has meaningfully improved the pursuit of quality of official statistics by allowing for more accurate population estimates through better deduplication of records.
 Before implementing this solution, Statistics Poland relied on user-defined rules for linking data without identifiers, which often resulted in imprecise population flow estimates due to missed matches or false positives.
@@
 2020/39/B/HS4/00941 (Towards census-like statistics for foreign-born populations – quality, data integration and estimation).
+This software was peer-reviewed and accepted by the pyOpenSci organisation (see https://github.com/pyOpenSci/software-submission/issues/232).
 CRediT authorship contribution statement
@@
 The datasets used in Section 4 are included with the BlockingPy package and are also publicly available from their original sources.
+Appendix A. Numerical comparison with blocklib
+We compare BlockingPy and blocklib in three synthetic datasets ($1,500$, $15,000$, and $150,000$ records) with duplicates $500$, $5,000$, and $50,000$, respectively.
+The results are averaged over 10 runs.
+The datasets were generated with geco3 and contain fictional personal information (given name, middle name, surname, municipality, date of birth, and nationality).
+CPU experiments used a 6-core Intel Core i5 with 16 GB RAM and Python 3.12.
+GPU runs used a NVIDIA RTX 3050 (4 GB VRAM) and Python 3.10.
+We report time (s), recall, reduction ratio (RR), and candidate pairs (millions).
+The comparison is also available in the documentation along with the code and data in the project repository.
+Preprocessing: for BlockingPy, we concatenate all fields of interest into a single text field; for blocklib, we tune algorithm parameters to approximately match the reduction ratio of the other methods to ease comparison.
+Table 4: CPU results on three datasets.
+Algorithm Size Time [s] Recall $R$ Pairs (M)
+P-Sig 1,500 0.051 (0.010) 0.599 (0.000) 0.996 (0.000) 0.004 (0.000)
+$\lambda$-fold LSH 1,500 0.193 (0.009) 0.466 (0.033) 0.991 (0.003) 0.010 (0.003)
+BlockingPy (voyager) 1,500 0.293 (0.029) 0.952 (0.006) 0.997 (0.000) 0.003 (0.000)
+BlockingPy (faiss_hnsw) 1,500 0.164 (0.049) 0.960 (0.000) 0.998 (0.000) 0.003 (0.000)
+BlockingPy (faiss_lsh) 1,500 0.185 (0.015) 0.955 (0.008) 0.997 (0.000) 0.003 (0.000)
+P-Sig 15,000 0.388 (0.087) 0.616 (0.000) 0.996 (0.000) 0.407 (0.000)
+$\lambda$-fold LSH 15,000 2.016 (0.251) 0.454 (0.028) 0.992 (0.003) 0.940 (0.337)
+BlockingPy (voyager) 15,000 6.777 (0.903) 0.876 (0.005) 1.000 (0.000) 0.041 (0.001)
+BlockingPy (faiss_hnsw) 15,000 12.029 (1.681) 0.913 (0.000) 1.000 (0.000) 0.031 (0.000)
+BlockingPy (faiss_lsh) 15,000 1.398 (0.148) 0.900 (0.004) 1.000 (0.000) 0.033 (0.001)
+P-Sig 150,000 3.444 (0.091) 0.609 (0.000) 0.996 (0.000) 40.231 (0.000)
+$\lambda$-fold LSH 150,000 19.848 (0.476) 0.450 (0.026) 0.992 (0.003) 95.064 (34.252)
+BlockingPy (voyager) 150,000 110.341 (3.815) 0.715 (0.005) 1.000 (0.000) 0.675 (0.021)
+BlockingPy (faiss_hnsw) 150,000 250.012 (4.232) 0.832 (0.001) 1.000 (0.000) 0.375 (0.001)
+BlockingPy (faiss_lsh) 150,000 56.544 (1.315) 0.818 (0.001) 1.000 (0.000) 0.400 (0.005)
+Table 5: GPU results on 150,000 records dataset.
+Algorithm Size Time [s] Recall $R$ Pairs (M)
+BlockingPy (gpu_faiss cagra) 150,000 51.084 (0.802) 0.827 (0.000) 1.000 (0.000) 0.380 (0.000)
+BlockingPy (gpu_faiss flat) 150,000 22.511 (0.508) 0.839 (0.000) 1.000 (0.000) 0.367 (0.000)
+BlockingPy (gpu_faiss ivf) 150,000 71.458 (2.185) 0.801 (0.004) 1.000 (0.000) 0.487 (0.021)

### v4

@@
 The objective of entity resolution is to link records without common unique identifiers that refer to the same entity (e.g., person, company).
-However, without identifiers, researchers need to specify which records to compare in order to calculate matching probability and reduce computational complexity.
-One solution is to deterministically block records based on some common variables, such as names, dates of birth or sex or use phonetic algorithms.
+However, without identifiers, researchers need to specify which records to compare in order to calculate matching probabilities and reduce computational complexity.
+One solution is to deterministically block records based on some common variables, such as names, dates of birth, or sex, or by using phonetic algorithms.
 However, this approach assumes that these variables are free of errors and completely observed, which is often not the case.
-To address this challenge, we have developed a Python package, BlockingPy, which uses blocking using modern approximate nearest neighbour search and graph algorithms to reduce the number of comparisons.
+To address this challenge, we have developed a Python package, BlockingPy, which relies on schema-agnostic blocking using modern approximate nearest neighbour search and graph algorithms to reduce the number of comparisons.
 The package supports both CPU and GPU execution.
@@
 C1 Current code version v0.2.8
-C2 Permanent link to code/repository used for this code version https://github.com/ncn-foreigners/BlockingPy
+C2 Permanent link to code/repository used for this code version N/A
 C3 Permanent link to Reproducible Capsule https://blockingpy.readthedocs.io/en/latest/examples/record_linkage.html
@@
 Existing deterministic blocking methods rely on user-defined exact-matching rules that assume key fields (e.g., name, date of birth) are error-free, potentially leading to missed links and false matches [4, 13, 14].
-Scientific community has developed several tools to address record linkage, including R packages such as RecordLinkage [6], reclin2 [7], fastLink [8], or Python implementations like splink [9], recordlinkage [10], Dedupe [11] or FEBRL [12].
+An alternative approach is schema-agnostic blocking [23, 31, 24], which concatenates all attribute values of a particular entity into a single textual representation before applying blocking (and can be combined with pre-trained language models).
+The scientific community has developed several tools to address record linkage, including R packages such as RecordLinkage [25], reclin2 [32], fastLink [11], or Python implementations like splink [16], recordlinkage [9], Dedupe [12], pyJedAI [21], or FEBRL [3].
 However, these solutions typically rely on user-defined deterministic rules that may miss certain groups of records or result in large blocks of records when data irregularities are present.
-Some solutions regarding these issues were proposed in the literature, for instance, the use of the SOUNDEX algorithm and other phonetic algorithms to block records on character strings that sound similar phonetically [13]; blocking using the Locality Sensitive Hashing (LSH; [14]), an approximate nearest neighbour (ANN) algorithm to search for similar records (see klsh package in R) or usage of Bloom filters, useful in particular for privacy-preserving record linkage (see, [15]; for implementation see blocklib in Python).
-The BlockingPy package uses state-of-the-art ANN search algorithms and graph-based indexing techniques, which allow it to work even with imperfect data.
+Some solutions to these issues were proposed in the literature, for instance, the use of the SOUNDEX algorithm and other phonetic algorithms to block records on character strings that sound similar phonetically [5]; blocking using the Locality Sensitive Hashing (LSH; [29]), an approximate nearest neighbour (ANN) algorithm to search for similar records (see, e.g., the klsh package in R) or the use of Bloom filters, useful in particular for privacy-preserving record linkage (see, [2]; for implementation see blocklib in Python).
+The BlockingPy package relies on the schema-agnostic approach and uses state-of-the-art ANN search algorithms and graph-based indexing techniques, which allow it to work even with imperfect data.
 BlockingPy implements an ANN-based blocking procedure ($n$-gram encoding or embeddings, neighbour search, graph components) behind a single pandas-style call, cutting custom rule-writing and setup time.
@@
 In Section 4, we provide two examples of probabilistic record linkage and deduplication.
-The paper finishes with a brief review of the possible impact and conclusions.
+The paper concludes with a brief review of the possible impact and conclusions.
 2 Existing software covering blocking procedures
-In this section, the focus is on Python, but readers are referred to the fastLink and reclin2 packages for more information on deterministic blocking procedures.
-The splink package allows for deterministic blocking rules using the function block_on() which specifies either the variables (e.g.
-block_on("city")) or more advanced combinations (e.g.
-block_on("substr(first_name, 1,1)", "surname")), which then are translated to a SQL query for the DuckDB (via the BlockingRuleCreator class).
+This section focuses on Python packages (splink, recordlinkage, and blocklib), but readers interested in deterministic blocking procedures are referred to the fastLink and reclin2 packages in R.
+For a comparison of Python packages, see Table 4 in the Appendix.
+The splink package allows for deterministic blocking rules using the function block_on() which specifies either the variables (e.g., block_on("city")) or more advanced combinations (e.g., block_on("substr(first_name, 1,1)", "surname")), which then are translated to a SQL query for the DuckDB (via the BlockingRuleCreator class).
 More advanced users may be interested in creating their own rules, which is possible through the CustomRule class.
 Splink allows users to specify the list of blocking rules for the Linker class.
-Splink allows users to inspect the blocking via the
-splink.blocking_analysis method.
+Splink allows users to inspect the blocking via the splink.blocking_analysis method.
 Additionally, splink supports blocking using phonetic algorithms such as Soundex, Metaphone, or Double Metaphone either by applying them directly within the blocking rules (e.g., block_on("soundex(surname)")), or by preprocessing columns to store phonetic encodings separately.
-In online documentation we show that using splink with BlockingPy together results in significantly improvement in the performance metrics by capturing comparison pairs that would otherwise be missed.
+In online documentation we show that using splink with BlockingPy together results in significant improvements in the performance metrics by capturing comparison pairs that would otherwise be missed.
 recordlinkage package supports deterministic blocking via the recordlinkage.Index class and its methods.
@@
 The only software that provides similar functionality is the blocklib package (cf.
-[16, 17]) and vicinity [18].
-The first uses probability signature and LSH-based Lambda Fold Redundant as described in [16].
+[34, 8]), vicinity [20], and pyJedAI.
+The first uses probability signature and LSH-based Lambda-Fold Redundant as described in [34].
 The second uses semhash [33], which allows deduplication of records using kNN-based embedding comparisons and offers exact-match blocking.
+The third provides an end-to-end entity resolution pipeline that includes multiple key-based blocking methods and comparison cleaning (meta-blocking) strategies, with an optional embedding-based candidate generation component based on FAISS.
 In contrast, BlockingPy allows various inputs as well as ANN algorithms and graphs, focusing on providing accurate blocking for probabilistic record linkage and deduplication.
-In Appendix Appendix A.
-Numerical comparison with blocklib, we compare BlockingPy with blocklib in terms of computational efficiency, accuracy, and number of pairs from the blocking procedures.
+In B, we compare BlockingPy with blocklib in terms of computational efficiency, accuracy, and number of pairs from the blocking procedures.
 While blocklib is faster, it results in a significantly higher number of missed pairs and several orders of magnitude more pairs.
-We also provide comparison with the GPU version of the package.
+We also provide a comparison with the GPU version of the package.
+In C we compare PyJedAI with BlockingPy on the Abt-Buy benchmark [15].
 3 Software description
@@
 Users can input either previously computed Document-Term Matrices (DTMs) or raw text data, which the package transforms into DTMs by constructing character $n$-gram representations of the input.
-Alternatively, users can choose to represent the input data as vector embeddings.
+Alternatively, users can represent the input data as vector embeddings.
+For models supported by the model2vec library, users can simply pass the model identifier as a string (e.g., "minishlab/potion-base-8m"), and the package computes the embeddings internally.
+For any other embedding model, users can precompute the embeddings separately and pass them to the package as NumPy arrays.
 Subsequently, similarity-based nearest neighbour search is performed for each record in either the input dataset (deduplication) or query dataset (record linkage).
@@
 Finally, BlockingPy groups records into blocks by identifying connected components in an undirected graph (using the igraph package [6]), generated by previous neighbourhood search results.
+In this graph, vertices represent records and edges connect pairs of records identified as neighbours.
+This step is necessary to handle transitive relationships (e.g., record A matches record B and record B matches record C).
 By using the block() method with the desired parameters on the Blocker instance, users can run the above-mentioned process to obtain blocking results.
@@
 Metric Formula Metric Formula
-Recall $\frac{TP}{TP+FN}$ Accuracy $\frac{TP+TN}{TP+TN+FP+FN}$
-Precision $\frac{TP}{TP+FP}$ Specificity $\frac{TN}{TN+FP}$
+Recall (Pairs Completeness) $\frac{TP}{TP+FN}$ Accuracy $\frac{TP+TN}{TP+TN+FP+FN}$
+Precision (Pairs Quality) $\frac{TP}{TP+FP}$ Specificity $\frac{TN}{TN+FP}$
 F1 Score $2\cdot\frac{\text{Precision}\times\text{Recall}}{\text{Precision}+\text{Recall}}$ False Positive Rate $\frac{FP}{FP+TN}$
@@
 1control_ann = {
-2 ’random_seed’: 2025
+2 ’random_seed’: 2025,
 3 ’hnsw’: {
@@
 Listing 3: Preprocessing the record linkage example data
-Before blocking, we merge all fields of interest (without ID) into a single field (i.e., txt) that we will pass to the algorithm as shown in Listing 4. This step is performed for both datasets.
+Before blocking, we merge all fields of interest (without ID) into a single field (i.e., txt) that we will pass to the algorithm as shown in Listing 4 (i.e, we use schema-agnostic blocking).
+This step is performed for both datasets.
 1for df in [census, cis]:
@@
 We also used this technique to determine the number of residents of Ukraine under temporary protection in Poland as of March 31, 2023 [28].
-The main motivation behind this was the problem of lack of identifiers as well as the transliteration from Ukrainian/Russian to Polish language (e.g.
+The main motivation behind this was the problem of the lack of identifiers as well as the transliteration from Ukrainian/Russian to Polish language (e.g.
 Олександр, should be Ołeksandr not Aleksandr).
@@
 We thank the reviewers for their comments on this manuscript.
-This software is a part of Tymoteusz Strojny’s Bachelor’s thesis at Poznań University of Economics and Business, Poland.
+This software was a part of Tymoteusz Strojny’s Bachelor’s thesis at Poznań University of Economics and Business, Poland.
 BlockingPy is based on the blocking package in R (see https://ncn-foreigners.github.io/blocking/).
@@
 The datasets used in Section 4 are included with the BlockingPy package and are also publicly available from their original sources.
-Appendix A. Numerical comparison with blocklib
+Appendix A Comparison of blocking methods implemented in splink, blocklib, recordLinkage, pyJedi and BlockingPy
+Table 4: Comparison of blocking methods implemented in selected software.
+Package Blocking methods
+splink Rule-based / deterministic blocking (SQL-style rules), phonetic blocking (e.g., Soundex, Metaphone).
+blocklib Probabilistic signature blocking (p-sig), LSH-based $\lambda$-fold redundant blocking (Bloom filter + LSH).
+recordlinkage Exact blocking, Sorted Neighbourhood, full indexing (all pairs), random indexing (sampled pairs).
+pyJedAI Token Blocking, Sorted Neighborhood, Extended Sorted Neighborhood, Q-Grams Blocking, Extended Q-Grams Blocking, Suffix Arrays Blocking, Extended Suffix Arrays Blocking, Embeddings-based NN blocking.
+BlockingPy ANN-based blocking via FAISS, HNSW (faiss / hnswlib / voyager), NN-Descent (pynndescent), Annoy, plus GPU indices (e.g., IVF, IVFPQ, CAGRA); blocks from connected components of the NN graph.
+Appendix B Numerical comparison with blocklib
 We compare BlockingPy and blocklib in three synthetic datasets ($1,500$, $15,000$, and $150,000$ records) with duplicates $500$, $5,000$, and $50,000$, respectively.
@@
 GPU runs used a NVIDIA RTX 3050 (4 GB VRAM) and Python 3.10.
-We report time (s), recall, reduction ratio (RR), and candidate pairs (millions).
+We report time (s), recall (Pairs Completeness), reduction ratio (RR), and candidate pairs (millions).
 The comparison is also available in the documentation along with the code and data in the project repository.
@@
 BlockingPy (gpu_faiss ivf) 150,000 71.458 (2.185) 0.801 (0.004) 1.000 (0.000) 0.487 (0.021)
+Appendix C Comparison with PyJedAI
+We compare BlockingPy and PyJedAI on the Abt-Buy clean-clean entity resolution benchmark, which consists of two product catalogs from the online retailers Abt.com and Buy.com.
+We use the Abt-Buy dataset as provided in the pyJedAI repository, which contains 1,076 records in each source and a gold-standard mapping between the two sources.
+The hardware setup is the same as in B; the experiments were run with Python 3.11 (with PyJedAI using NumPy 2.4.1 and BlockingPy using NumPy 1.26.4).
+We report runtime (s), recall, precision, $F_{1}$, and the number of candidate comparison pairs.
+For BlockingPy, we use the HNSW backend with the minishlab/potion-base-32M embedding model; all remaining parameters are left at their defaults.
+For PyJedAI, we use the sminilm vectorizer with $k=5$ nearest neighbors and Euclidean distance.
+The data (and an example workflow) are available in the pyJedAI documentation/repository.
+The BlockingPy workflow used for this comparison is likewise available in the BlockingPy repository.
+Table 7: Comparison on the Abt-Buy benchmark
+Algorithm Time [s] Recall Precision $F_{1}$ Pairs PyJedAI 52.6 0.9377 0.1875 0.3126 5380 BlockingPy 4.4 0.8234 0.8234 0.8234 1076
+Note: The identical recall, precision, and $F_{1}$ values reported here are expected in this setting and are not a reporting error: in BlockingPy we use $k=1$ (one candidate per record) and, for this experiment, the graph-based step did not find additional links, so evaluation is effectively performed on the same one-to-one pair set; under these conditions these three metrics coincide.